OpenStudy (anonymous):

What is a polynomial function in standard form with zeroes 1, 2, –3, and –3?

OpenStudy (anonymous):

zero at 1 means (x-1) can you figure the others?

OpenStudy (anonymous):

yeah but what would it be? g(x) = x4 + 3x3 –7x2 – 15x + 18 g(x) = x4 + 3x3 – 7x2 + 2x + 18 g(x) = x4 – 3x3 + 7x2 + 15x + 18 g(x) = x4 – 3x3 – 7x2 + 15x + 18

OpenStudy (anonymous):

write the other 3 then multiply everything and and you should get it